# Adaptive Progressive Attention Graph Neural Network for EEG Emotion Recognition

### v1

@@ Abstract
In recent years, numerous neuroscientific studies have shown that human emotions are closely linked to specific brain regions, with these regions exhibiting variability across individuals and emotional states. To fully leverage these neural patterns, we propose an Adaptive Progressive Attention Graph Neural Network (APAGNN), which dynamically captures the spatial relationships among brain regions during emotional processing. The APAGNN employs three specialized experts that progressively analyze brain topology. The first expert captures global brain patterns, the second focuses on region-specific features, and the third examines emotion-related channels. This hierarchical approach enables increasingly refined analysis of neural activity. Additionally, a weight generator integrates the outputs of all three experts, balancing their contributions to produce the final predictive label. Extensive experiments on three publicly available datasets (SEED, SEED-IV and MPED) demonstrate that the proposed method enhances EEG emotion recognition performance, achieving superior results compared to baseline methods.
@@
Progressive attention graph neural network, electroencephalography (EEG), EEG emotion recognition
@@ I Introduction
Emotions significantly impact various aspects of human daily life and mental health, influencing our decision-making, motivation, attention, memory, problem-solving capabilities[1, 2]. Consequently, it is necessary to develop methods to objectively and accurately identify human emotions. As the key step to make machines understand human emotions, emotion recognition has received substantial attention from human–machine interaction (HMI) and pattern recognition communities in recent years[3, 4, 5, 6]. Most of these studies primarily utilize two typical signals, i.e., external and internal responses. The external responses mainly include some behavioral data, such as facial expression[7], speech signals[8], conversational data on social media platforms[9]. The internal responses are based on physiological signals, such as electromyography (EMG)[10], electrocardiogram (ECG)[11], and electroencephalogram (EEG)[12]. Neuroscience research indicates that physiological signals are closer to the source of emotion than behavioral signals. Thus, more and more researchers focus on this field during the past several years.
@@
EEG is a widely used technique for recording the electrophysiological activity of neurons in the cerebral cortex via electrodes placed on the scalp [13]. As a typical physiological signal, EEG has shown considerable promise for decoding human emotions  [14, 15, 16, 17]. Over the past decades, numerous methods have been proposed to effectively analyze EEG emotional signals. Convolutional neural networks (CNNs), in particular, have been widely used for EEG emotion recognition. For example, Hasan et al. [18] employed fast Fourier transform (FFT) and CNNs to classify 64 emotions, achieving notable accuracy in valence and arousal dimensions. Similarly, Ali et al. [19] utilized a Capsule Network (CapsNet) that outperformed traditional support vector machines (SVMs) and CNNs, achieving average accuracies of 80.22% and 85.41%, respectively. Ahmad et al. [20] focused on a CNN-based model for classifying emotions into positive, neutral, and negative categories, attaining over 93% accuracy across all categories. Recent advancements include He et al.’s [21] CNN architecture for complex EEG signal classification and Aldawsari et al.’s [22] optimized 1D-CNN-based processh, which significantly improved efficiency. Hybrid models combining CNNs with other techniques, such as LSTM and ensemble learning, have also shown promising results. For instance, Ali et al. [23] and Yuvaraj et al. [24] demonstrated high accuracies, with Ali’s model reaching an impressive 98%. Other notable innovations include the integration of differential entropy with CNN-BiLSTM by Cui et al. [25], which achieved over 94% accuracy on the DEAP and SEED datasets, and the introduction of CIT-EmotionNet by Lu et al. [26], which combined CNN and Transformer models to outperform previous methods. Huang et al. [27] incorporated attention mechanisms within a CNN-BiLSTM framework, achieving near-perfect accuracies in multi-class tasks. Further studies, such as those by Saha et al. [28], Li et al. [29], and Tao et al. [30], explored wavelet decomposition, multi-scale CNNs, and attention mechanisms, all contributing to improved emotion recognition performance from EEG signals. Wang et al. [3] introduced self-supervised learning to CNNs, enhancing both resource efficiency and performance. Recent approaches include Jin et al.’s [31] CNN-Transformer network for fNIRS-based emotion recognition and Farokhah et al.’s [32] simplified 2D CNN with channel selection to improve inter-subject accuracy. Asif et al. [33] demonstrated the potential of combining various convolution layers within a CNN framework for subject-independent tasks.
@@
While these algorithms have shown impressive performance, they often overlook the inter-channel topological relationships inherent in EEG channel spatial features. This limitation arises from CNNs’ inability to process non-Euclidean spaces, such as graphs and manifolds. To better capture the spatial features of EEG signals, several Graph Neural Network (GNN)-based approaches have been proposed, yielding promising results. Zheng et al. [34] introduced a Hierarchy Graph Convolution Network (ERHGCN) that achieved accuracies of 90.56% and 88.79% for valence and arousal, respectively, demonstrating the potential of hierarchical GCN structures for emotion recognition. Saboksayr et al. [35] applied Graph Signal Processing (GSP) techniques, showing superior performance compared to traditional methods and highlighting the efficacy of graph-based approaches. Gilakjani et al. [36] combined GNNs with contrastive learning and GAN-based data augmentation, leading to enhanced classification performance on the DEAP and MAHNOB-HCI datasets. Li et al. [29] developed a model that learns discriminative graph topologies in EEG networks, achieving an average accuracy of 84.56% in online experiments, which is particularly important for affective brain-computer interfaces. Klepl et al. [37] provided a comprehensive survey on GNN applications in EEG signal classification, providing valuable insights into their strengths and limitations. In the realm of deep learning, Abdulrahman et al. [38] achieved notable accuracies of 70.89% in binary-class and 90.33% in multi-class emotion recognition tasks using advanced deep learning models. DGCNN, proposed by Song et al. [39], dynamically learns the adjacency matrix of the graph to establish relationships among EEG channels. Zhong et al.’s RGNN [13] enhances the robustness of GNN models against cross-subject variations and noisy labels by incorporating two regularization techniques. The IAG model by Song et al. [40] adaptively generates directed graph connections from input graphs, allowing for exploration of intrinsic relationships between EEG regions.
@@
Despite significant advancements in EEG-based emotion recognition, current methods still face challenges in dynamically capturing the complex relationships between emotional patterns and brain functional regions. One major challenge is the significant variation in emotion-related brain activity between individuals, which demands more adaptive recognition approaches. To address these challenges, we propose a novel Adaptive Progressive Attention Graph Neural Network (APAGNN) that finely screen out critical EEG channels across different subjects through three specialized experts. Specifically, three experts work collaboratively through a progressive analysis pipeline. The first expert analyzes global brain topology patterns, followed by the second expert that identifies emotion-specific regions, while the third expert focuses on critical channels. This hierarchical refinement process enables a comprehensive understanding of emotion-related neural patterns at multiple levels of granularity. The APAGNN model incorporates several innovative components to enhance its effectiveness. To encourage diversity in feature learning while preventing redundancy among experts, we develop a diversity-preserving training strategy that maximizes Jensen-Shannon (JS) Divergence among expert probability distributions. Additionally, we design a dynamic expert fusion method that optimally integrates the outputs from multiple experts for final classification.
@@
To the best of our knowledge, this is the first work to exploit the discrimination for emotional EEG expression from global to region brain. Our experimental results demonstrate the effectiveness of this progressive attention model. Besides, we also investigate the effect of the number of experts on emotion recognition performance, compare dynamic and static attention mechanisms, and analyze how these factors influence the overall accuracy of emotion classification.
@@
Figure 1: The overall architecture of APAGNN. In the multi-expert architecture, each expert learns discriminative features for emotion classification while generating attention maps to guide subsequent experts. The first two experts progressively refine the graph structure through attention operations and node pruning, optimizing the input for the next expert in the sequence. The dynamic expert fusion module then integrates the experts’ predictions using a weight generator to produce the final model output.
@@ II-A Graph Neural Network
An undirected and connected graph is represented as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ denotes the set of vertices (or nodes) and $\mathcal{E}$ denotes the set of edges, which are unordered pairs of vertices. For a graph with $|\mathcal{V}|=N$ nodes, its structure can be characterized by an adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$, where $\mathbf{A}_{ij}=1$ if vertices $i$ and $j$ are connected by an edge, and $\mathbf{A}_{ij}=0$ otherwise.
@@
To address the challenges of localized filtering and computational complexity, Defferrard et al. [41] proposed the use of Chebyshev polynomials to approximate the spectral filters. This approach leverages the properties of Chebyshev polynomials to construct an efficient filtering process by considering only a limited range of eigenvalues. Consequently, the convolution operation becomes more efficient, enabling the practical application of GCNs to larger graphs while maintaining the ability to capture localized patterns in the data.
@@ II-B Attention Mechanism
The attention mechanism in deep learning draws inspiration from the selective attention process observed in human vision. When processing visual information, the human visual system focuses on salient regions that are pertinent for decision-making while disregarding less relevant portions [42]. Similarly, in deep learning applications, certain parts of the input may hold more significance for decision-making than others. In recent years, various attention mechanisms have been proposed to capture this selective processing, including Class Activation Mapping (CAM), Gradient-weighted Class Activation Mapping (Grad-CAM), Saliency Maps, and both hard and soft attention strategies. The integration of these attention modules into neural networks has demonstrated significant performance improvements across numerous studies [43, 44, 45, 46].
In computer vision tasks, feature maps from the final convolutional layer are known to encode rich semantic information. These feature maps are weighted, summed, and then upscaled to the original image size to obtain the CAM for a specific category. The resulting attention map highlights regions of the input image that are most influential for classifying a specific category. Grad-CAM, proposed by Selvaraju et al. [47], extends the original CAM approach by leveraging class-specific gradient information as weights for the feature maps. Thus, Grad-CAM can be applied to a wider range of network architectures without the need for architectural modifications. The Grad-CAM computation is formally expressed as $L_{\text{Grad-CAM}}^{c}=\operatorname{ReLU}\left(\sum_{k}\alpha_{k}^{c}F_{k}\right)$, where $L_{\text{Grad-CAM}}^{c}$ represents the Grad-CAM heatmap for class $c$, and $F_{k}$ denotes the $k$-th feature map from the final convolutional layer. The ReLU activation ensures that only features positively contributing to the class prediction are preserved in the attention map. The weight coefficient $\alpha_{k}^{c}$ for the $k$-th feature map is calculated by backpropagating the gradients of the class score $y^{c}$ with respect to the feature map $F_{k}$: $\alpha_{k}^{c}=\frac{1}{WH}\sum_{i=0}^{W}\sum_{j=0}^{H}\frac{\partial y^{c}}{\partial F_{k}(i,j)}$, where $W$ and $H$ denote the width and height of feature map $F_{k}$, respectively, and $F_{k}(i,j)$ represents the activation value of the $k$-th feature map at spatial location $(i,j)$. The score $y^{c}$ corresponds to the logit or probability for class $c$ from the network’s output. The weighting coefficient $\alpha_{k}^{c}$ signifies the importance of the $k$-th feature map $F_{k}$ concerning the target class $c$, effectively measuring its contribution to the prediction score $y^{c}$.
@@ III Methodology
To specify the proposed method clearly, we illustrate the framework of the APAGNN model in Fig. 1. Its goal is to capture the more discriminative EEG representation for emotion recognition. We adopt three steps to achieve this goal. The first step is to build the spatial relationships among brain regions from global to local regions progressively. Subsequently, to encourage diversity in feature learning while preventing redundancy among experts, we implement a knowledge diversity-preserving operation. Third, a dynamic expert fusion strategy is designed to integrate the knowledge from multiple experts for final emotion prediction. The overall process is described as follows.
@@ III-A Multi-Expert Progressive Learning
To model the complex spatial relationships among different neural regions at multiple scales, we employ three specialized experts that progressively analyzes brain activity from global patterns to regional features and ultimately to electrode-specific characteristics. Each expert focuses on a different spatial scale, allowing for both broad and fine-grained feature extraction. The structure of each expert is presented in Fig. 2. Each expert performs two primary tasks: learning discriminative features for emotion classification and generating attention maps to transfer knowledge. By utilizing three experts with this structure, we achieve the goal of progressive learning.
Graph-based EEG representation Learning. To effectively capture the spatial dependencies in EEG signals, we leverage GNNs with Chebyshev filters to construct each expert. The process begins by transforming raw EEG samples into graph representations, where each EEG channel serves as a node. Specifically, we first extract differential entropy (DE) features in delta (1-4 Hz), theta (4-8 Hz), alpha (8–13 Hz), beta (13–30 Hz), and gamma (36–44 Hz) frequency bands. These features serve as the initial node attributes, resulting in a feature matrix $\mathbf{X}\in\mathbb{R}^{C\times F}$, where $C$ represents the number of electrodes, and $F$ denotes the number of frequency bands. To model the spatial relationships between electrodes, we construct an adjacency matrix $\mathbf{A}\in\mathbb{R}^{C\times C}$ based on the physical arrangement of the electrodes [48]. Through this process, each raw EEG sample is represented as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})=(\mathbf{X},\mathbf{A})$, where $\mathcal{V}$ and $\mathcal{E}$ represent the sets of nodes and edges in the graph, respectively. The graph convolution operation, implemented using Chebyshev polynomial formulation, is defined as:
$\mathbf{y}=\sum_{k=0}^{K-1}\theta_{k}T_{k}(\widetilde{\mathbf{L}})\,\mathbf{X}_{,}$ (1)
where $\mathbf{\theta}_{k}$ represents learnable parameters for the $k$-th order Chebyshev polynomial, and $T_{k}(\widetilde{\mathbf{L}})$ denotes the $k$-th order Chebyshev polynomial evaluated at the scaled Laplacian $\widetilde{\mathbf{L}}$. To efficiently compute higher-order Chebyshev polynomials, we use the recurrence relation. Define $\bar{\mathbf{X}}_{k}=T_{k}(\widetilde{\mathbf{L}})\mathbf{X}$, with $\bar{\mathbf{X}}_{0}=\mathbf{X}$, $\bar{\mathbf{X}}_{1}=\tilde{\mathbf{L}}\mathbf{X}$, and $\bar{\mathbf{X}}_{k}=2\widetilde{\mathbf{L}}\bar{\mathbf{X}}_{k-1}-\bar{\mathbf{X}}_{k-2}$. Thus, the graph convolution operation can be expressed as:
@@
For a single Chebyshev filter, the learnable coefficients are represented by $\Theta_{i}=[\theta_{0},\theta_{1},...,\theta_{K-1}]^{T}\in\mathbb{R}^{K\times 1}$. To learn diverse feature transformations, we implement $D$ parallel filters, each learning a different transformation of the node features. The complete set of learnable parameters is denoted as $\Theta=[\Theta_{0},\Theta_{1},...,\Theta_{D-1}]\in\mathbb{R}^{K\times D}$. After computing the outputs for each node and filter across all $K$ Chebyshev orders, we concatenate the results and apply the weights $\Theta$. Finally, through the ReLU activation function, the learned EEG representation $\mathbf{H}_{i}\in\mathbb{R}^{C\times D}$ for the $i$-th expert are obtained:
@@
The learned features $\mathbf{H}_{i}$ is then flattened to feed into a fully connected layer to obtain class probability scores $\mathbf{S}_{i}\in\mathbb{R}^{E}$ for each emotion class, where $E$ represents the number of emotion classes. The classification loss for the $i$-th expert is computed as:
@@
where $y_{i}^{l}$ is the ground truth label (one-hot encoded) for class $l$, and $S_{i}^{l}$ represents the predicted probability for class $l$ by the $i$-th expert.
@@
Specifically, for a target emotion class index $l$, we compute the gradient of the predicted score $S_{i}^{l}$ with respect to the feature maps $\mathbf{H}_{i}$ through backpropagation. The gradients are processed through global average pooling to derive the importance score $\alpha_{d}^{l}$ for each feature dimension $d$:
@@
where $C$ denotes the total number of EEG channels. This gradient-based approach reflects the importance of each feature dimension on the decision.
Next, the weights $\alpha_{d}^{l}$ and the feature map $\mathbf{H}_{i}$ are linearly combined to generate the attention map for class $l$. The attention weights $\alpha_{d}^{l}$ are averaged across feature dimensions to obtain the final importance score $\mathbf{I}_{i}$ for each EEG channel:
$\mathbf{I}_{i}=\operatorname{ReLU}\left(\sum_{d=0}^{D-1}\mathbf{H}_{i}\right),$ (6)
@@
Expert Progressive Learning. The three experts work in sequence to progressively refine the analysis of EEG signals. The pipeline begins with the first expert conducting a global-scale analysis on the original graph $\mathcal{G}$, generating both feature representation $\mathbf{H}_{1}$ and attention map $\mathbf{\Phi}_{1}$ that identifies emotionally salient regions. Building upon this foundation, the second expert receives both the original graph $\mathcal{G}$ and the attention map $\mathbf{\Phi}_{1}$ from the first expert to perform region-scale analysis. By utilizing $\mathbf{\Phi}_{1}$ to mask irrelevant nodes in $\mathcal{G}$, it focuses on the emotion-relevant areas. The second expert then applies the same graph convolution and attention computation process with its own learnable parameters, producing a refined feature representation $\mathbf{H}_{2}$ and an attention map $\mathbf{\Phi}_{2}$ with important electrodes. The final stage of analysis is performed by the third expert, which combines $\mathcal{G}$ and $\mathbf{\Phi}_{2}$ to conduct electrode-scale analysis, yielding the feature representation $\mathbf{H}_{3}$ that captures fine-grained spatial patterns in the EEG signals.

### v2

@@ Abstract
In recent years, numerous neuroscientific studies demonstrate that specific areas of the brain are connected to human emotional responses, with these regions exhibiting variability across individuals and emotional states. To fully leverage these neural patterns, we propose an Adaptive Progressive Attention Graph Neural Network (APAGNN), which dynamically captures the spatial relationships among brain regions during emotional processing. The APAGNN employs three specialized experts that progressively analyze brain topology. The first expert captures global brain patterns, the second focuses on region-specific features, and the third examines emotion-related channels. This hierarchical approach enables increasingly refined analysis of neural activity. Additionally, a weight generator integrates the outputs of all three experts, balancing their contributions to produce the final predictive label. Extensive experiments conducted on SEED, SEED-IV and MPED datasets indicate that our method enhances EEG emotion recognition performance, achieving superior results compared to baseline methods.
@@
Progressive attention, electroencephalography (EEG), graph neural network, EEG emotion recognition.
@@ I Introduction
Emotions significantly impact various aspects of human daily life and psychological health, influencing our decision-making, motivation, attention, memory, problem-solving abilities [1, 2]. Consequently, it is necessary to develop methods to objectively and accurately identify human emotions. The ability for machines to understand human feelings has become an important field of study, drawing significant attention from Human–Machine Interaction (HMI) and pattern recognition researchers lately [3, 4, 5, 6]. Most of these studies primarily utilize two typical signals, i.e., external and internal responses. External responses mainly include some behavioral data, such as facial expression[7], speech signals[8], conversational data on social media platforms[9]. Internal responses are based on physiological signals, including electromyography (EMG)[10], electrocardiogram (ECG)[11], and electroencephalogram (EEG)[12]. Neuroscience research suggests that physiological signals provide more direct access to emotional origins than behavioral indicators. Thus, more and more researchers focus on this field during the past several years.
@@
EEG is a widely used technique for recording the electrophysiological activity of neurons in the cerebral cortex via electrodes attached to the scalp [13]. As a typical physiological signal, EEG has shown considerable promise in decoding human emotions [14, 15, 16, 17]. Over the past decades, numerous approaches have emerged for efficiently decoding emotions from EEG signals. Among various methods, convolutional neural networks (CNNs) are extensively employed in recognizing emotions from EEG data. For example, Hasan et al. [18] utilize Fast Fourier Transform (FFT) and CNNs to classify 64 emotions, achieving notable accuracy in valence and arousal dimensions. Similarly, Ali et al. [19] utilize a Capsule Network (CapsNet) that outperforms traditional support vector machines (SVMs) and CNNs, achieving average accuracies of 80.22% and 85.41%, respectively. Ahmad et al. [20] focused on a CNN-based model for classifying emotions into positive, neutral, and negative categories, attaining over 93% accuracy across all categories. Recent advancements include He et al.’s [21] CNN architecture for complex EEG signal classification and Aldawsari et al.’s [22] optimized 1D-CNN-based process, which significantly improve efficiency. Hybrid models combining CNNs with other techniques, such as LSTM and ensemble learning, have also shown promising results. As an example, Ali et al. [23] and Yuvaraj et al. [24] demonstrate high accuracies, with Ali’s model reaching an impressive 98%. Other notable innovations include the integration of differential entropy with CNN-BiLSTM by Cui et al. [25], which achieves over 94% accuracy on DEAP and SEED datasets, and the introduction of CIT-EmotionNet by Lu et al. [26], which combines CNN and Transformer models to outperform previous methods. Huang et al. [27] incorporate attention mechanisms within a CNN-BiLSTM framework, achieving near-perfect accuracies in multi-class tasks. Further studies, such as those by Saha et al. [28], Li et al. [29], and Tao et al. [30], explore wavelet decomposition, multi-scale CNNs, and attention mechanisms, all contributing to improved emotion recognition performance from EEG signals. Wang et al. [3] introduces self-supervised learning to CNNs, enhancing both resource utilization and performance. Recent approaches include Jin et al.’s [31] CNN-Transformer network for fNIRS-based emotion recognition and Farokhah et al.’s [32] simplified 2D CNN with selective channel to improve inter-subject accuracy. Asif et al. [33] demonstrate the potential of combining various convolution layers within a CNN framework for subject-independent tasks.
@@
While these algorithms have shown impressive performance, they often overlook the inter-channel topological relationships inherent in EEG channel spatial features. This limitation arises from CNNs’ inability to process non-Euclidean spaces, such as graphs and manifolds. To better capture the spatial relationships within EEG signals, several Graph Neural Network (GNN)-based approaches have been proposed, yielding promising results. Zheng et al. [34] propose a Hierarchy Graph Convolution Network (ERHGCN) that achieves classification performance of 90.56% for valence dimension and 88.79% for arousal dimension, demonstrating the potential of hierarchical GCN structures for emotion recognition. Saboksayr et al. [35] apply Graph Signal Processing (GSP) techniques, showing enhanced performance when compared with traditional methods and highlighting the efficacy of graph-based approaches. Gilakjani et al. [36] combine GNNs with contrastive learning and GAN-based data augmentation, leading to enhanced classification performance on both DEAP and MAHNOB-HCI datasets. Li et al. [29] develop a model that learns discriminative graph topologies in EEG networks, achieving an average accuracy of 84.56% in online experiments, which is particularly important for affective brain-computer interfaces. Klepl et al. [37] present a comprehensive survey on GNN applications in EEG signal classification, providing valuable insights into their advantages and limitations. In the realm of deep learning, Abdulrahman et al. [38] achieve notable accuracies of 70.89% in binary classification and 90.33% for multi-class emotion recognition tasks using advanced deep learning models. DGCNN, proposed by Song et al. [39], dynamically learns the graph’s adjacency matrix to establish spatial relationships. Zhong et al.’s RGNN [13] enhances GNN model robustness against cross-subject variations and noisy labels by incorporating two regularization techniques. The IAG model by Song et al. [40] adaptively generates directed graph connections from input graphs, allowing for exploration of intrinsic relationships between EEG regions.
@@
Although considerable progress has been made in emotion recognition using EEG, current approaches still face challenges in dynamically capturing the complex relationships between emotional patterns and brain functional regions. One major challenge is the significant variation in emotion-related brain activity between individuals, which demands more adaptive recognition approaches. To address these challenges, we propose a novel Adaptive Progressive Attention Graph Neural Network (APAGNN) that finely screens out critical EEG channels across different subjects through three specialized experts. Specifically, three experts work collaboratively through a progressive analysis pipeline. The first expert analyzes global brain topology patterns, followed by the second expert that identifies emotion-specific regions, while the third expert focuses on critical channels. This hierarchical refinement process enables a comprehensive understanding of emotion-related neural patterns at multiple levels of granularity. The APAGNN model incorporates several innovative components to enhance its effectiveness. To encourage diversity in feature learning while preventing redundancy among experts, we develop a diversity-preserving training strategy that maximizes Jensen-Shannon (JS) Divergence among expert probability distributions. Additionally, we design a dynamic expert fusion method that optimally integrates the outputs from multiple experts for final classification.
@@
To our knowledge, this is the first work to exploit the discrimination for emotional EEG expression from global to region brain. Our experimental results highlight the effectiveness of this progressive attention model. Besides, we also investigate the impact that varying numbers of experts have on the performance of emotion recognition, compare dynamic and static attention mechanisms, and analyze how these factors influence the overall accuracy of emotion classification.
@@
Figure 1: The architecture of APAGNN. In the multi-expert architecture, each expert learns discriminative features for emotion classification while generating attention maps to guide subsequent experts. The first two experts progressively refine the graph structure through attention operations and node pruning, optimizing the input for the next expert in the sequence. The dynamic expert fusion module then integrates the experts’ predictions using a weight generator to produce the final model output.
@@ II-A Graph Neural Network
We represent an undirected and connected graph as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of vertices (or nodes) and $\mathcal{E}$ corresponds to the set of edges, which are unordered pairs of vertices. For a graph with $|\mathcal{V}|=N$ nodes, its structure can be characterized by the adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$, where $\mathbf{A}_{ij}=1$ if a connection exists between vertices $i$ and $j$, and $\mathbf{A}_{ij}=0$ otherwise.
@@
To address the challenges of localized filtering and computational complexity, Defferrard et al. [41] introduced Chebyshev polynomials to approximate the spectral filters. This approach leverages the properties of Chebyshev polynomials to construct an efficient filtering process by considering only a limited range of eigenvalues. Consequently, the convolution operation becomes more efficient, enabling the practical application of GCNs to larger graphs while maintaining the ability to capture localized characteristics.
@@ II-B Attention Mechanism
The concept of attention in deep learning was inspired by human visual attention behaviors. During visual processing, humans concentrate on salient regions that are pertinent for decision-making while disregarding less relevant portions [42]. Similarly, in deep learning applications, certain parts of the input may hold more significance for decision-making than others. Over the past few years, various attention-based approaches have been designed to capture this selective processing, including Class Activation Mapping (CAM), Gradient-weighted Class Activation Mapping (Grad-CAM), Saliency Maps, and both hard and soft attention strategies. The integration of these attention modules into neural networks has demonstrated significant performance improvements across numerous studies [43, 44, 45, 46].
In computer vision tasks, feature maps from the final convolutional layer are known to encode rich semantic information. These feature maps are weighted, summed, and then upscaled to the original image size to obtain the CAM for a specific category. The resulting attention map highlights regions of the input that are most influential on classification decisions. Grad-CAM, proposed by Selvaraju et al. [47], extends the original CAM approach by leveraging class-specific gradient information as weights for the feature maps. Thus, Grad-CAM works with many different types of network structures without the need for architectural modifications. The Grad-CAM computation is formally expressed as $L_{\text{Grad-CAM}}^{c}=\operatorname{ReLU}\left(\sum_{k}\alpha_{k}^{c}F_{k}\right)$, where $L_{\text{Grad-CAM}}^{c}$ represents the corresponding Grad-CAM for category $c$, and $F_{k}$ denotes the $k$-th feature map from the final convolutional layer. The ReLU activation ensures that only features positively contributing to the class prediction are preserved in the attention map. The weight coefficient $\alpha_{k}^{c}$ for the $k$-th feature map is calculated by backpropagating the gradients of the specific class score $y^{c}$ with respect to the feature map $F_{k}$: $\alpha_{k}^{c}=\frac{1}{WH}\sum_{i=0}^{W}\sum_{j=0}^{H}\frac{\partial y^{c}}{\partial F_{k}(i,j)}$, where $W$ and $H$ are the width and height of feature map $F_{k}$, respectively, and $F_{k}(i,j)$ represents the activation value of the $k$-th feature map at spatial location $(i,j)$. The score $y^{c}$ corresponds to the logit or probability for class $c$ from the network’s output. The weighting coefficient $\alpha_{k}^{c}$ signifies the importance of the $k$-th feature map $F_{k}$ concerning the target class $c$, effectively measuring its contribution to the prediction score $y^{c}$.
@@ III Methodology
For clarity regarding our methodology, Fig. 1 presents the APAGNN model structure. Its goal is to capture more discriminative EEG representation for emotion recognition. We employ three steps to attain this objective. The first step focuses on building the spatial relationships among brain regions from global to local regions progressively. Subsequently, to encourage diversity in feature learning while preventing redundancy among experts, we implement a knowledge diversity-preserving operation. Third, a dynamic expert fusion strategy is designed to integrate the knowledge from multiple experts for final emotion prediction. We detail the complete procedure below.
@@ III-A Multi-Expert Progressive Learning
To model the complex spatial relationships among different neural regions at multiple scales, we employ three specialized experts that progressively analyze brain activity from global patterns to regional features and ultimately to electrode-specific characteristics. Each expert focuses on a different spatial scale, allowing for both broad and fine-grained feature extraction. The structure of each expert is presented in Fig. 2. Each expert performs two primary tasks: learning discriminative features for emotion classification and generating attention maps to transfer knowledge. By utilizing three experts with this structure, we achieve the goal of progressive learning.
Graph-based EEG representation Learning. To effectively capture the spatial dependencies in EEG signals, we leverage GNNs with Chebyshev filters to construct each expert. The process begins by transforming raw EEG samples into graph representations, where each EEG electrode serves as a vertex. Specifically, we first extract differential entropy (DE) features from five frequency bands. These features serve as the initial node attributes, yielding a feature matrix $\mathbf{X}\in\mathbb{R}^{C\times F}$, in which $C$ represents the electrode count, and $F$ indicates the frequency band count. To model the spatial relationships between electrodes, we construct an adjacency matrix $\mathbf{A}\in\mathbb{R}^{C\times C}$ based on the physical arrangement of the electrodes [48]. Through this process, each raw EEG sample is converted into a graph structure $\mathcal{G}=(\mathcal{V},\mathcal{E})=(\mathbf{X},\mathbf{A})$, with $\mathcal{G}$ denoting a graph with vertices $\mathcal{V}$ and edges $\mathcal{E}$. The graph convolution operation, implemented using Chebyshev polynomial formulation, is given by:
$\mathbf{y}=\sum_{k=0}^{K-1}\theta_{k}T_{k}(\widetilde{\mathbf{L}})\,\mathbf{X},$ (1)
where $\mathbf{\theta}_{k}$ represents learnable parameters for the $k$-th order Chebyshev polynomial, and $T_{k}(\widetilde{\mathbf{L}})$ denotes the $k$-th order Chebyshev polynomial evaluated at the scaled Laplacian $\widetilde{\mathbf{L}}$. To efficiently compute higher-order Chebyshev polynomials, we use the recurrence relation. Define $\bar{\mathbf{X}}_{k}=T_{k}(\widetilde{\mathbf{L}})\mathbf{X}$, with $\bar{\mathbf{X}}_{0}=\mathbf{X}$, $\bar{\mathbf{X}}_{1}=\tilde{\mathbf{L}}\mathbf{X}$, and $\bar{\mathbf{X}}_{k}=2\widetilde{\mathbf{L}}\bar{\mathbf{X}}_{k-1}-\bar{\mathbf{X}}_{k-2}$. Thus, we can formulate the convolution operation on graphs as:
@@
For a single Chebyshev filter, the learnable coefficients are represented by $\Theta_{i}=[\theta_{0},\theta_{1},...,\theta_{K-1}]^{T}\in\mathbb{R}^{K\times 1}$. To learn diverse feature transformations, we implement $D$ parallel filters, each learning a different transformation of the node features. The complete set of learnable parameters is denoted as $\Theta=[\Theta_{0},\Theta_{1},...,\Theta_{D-1}]\in\mathbb{R}^{K\times D}$. After computing the outputs for each node and filter across all $K$ Chebyshev orders, we concatenate the results and apply the weights $\Theta$. Finally, through the ReLU activation function, the learned EEG representation $\mathbf{H}_{i}\in\mathbb{R}^{C\times D}$ for the $i$-th expert is obtained:
@@
The learned features $\mathbf{H}_{i}$ are then flattened to input to a fully connected layer to obtain class probability scores $\mathbf{S}_{i}\in\mathbb{R}^{E}$ for each emotion class, with $E$ indicating the total emotion categories. The $i$-th expert’s classification loss is calculated using:
@@
where $y_{i}^{l}$ represents the one-hot encoded actual label for class $l$, and $S_{i}^{l}$ represents the probability of class $l$ predicted by the $i$-th expert.
@@
Specifically, for a target emotion class index $l$, we calculate the gradient between predicted probability $S_{i}^{l}$ and feature representations $\mathbf{H}_{i}$ through backpropagation. The gradients are processed through global average pooling to derive the importance score $\alpha_{d}^{l}$ for each feature dimension $d$:
@@
This gradient-based approach reflects the significance of each feature dimension on the classification decision. Next, the weights $\alpha_{d}^{l}$ and the feature map $\mathbf{H}_{i}$ are linearly combined to generate the attention map for class $l$. By taking the mean of attention weights $\alpha_{d}^{l}$ across feature dimensions, we derive the final channel importance $\mathbf{I}_{i}$:
$\mathbf{I}_{i}=\operatorname{ReLU}\left(\sum_{d=0}^{D-1}\alpha_{d}^{l}\mathbf{H}_{i}\right),$ (6)
@@
Expert Progressive Learning. The three experts work in sequence to progressively refine the analysis of EEG signals. The pipeline begins with the first expert conducting a global-scale analysis on the original graph $\mathcal{G}$, generating both feature representation $\mathbf{H}_{1}$ and attention map $\mathbf{\Phi}_{1}$ that identifies emotionally salient regions. Building upon this foundation, the second expert receives both the original graph $\mathcal{G}$ and the attention map $\mathbf{\Phi}_{1}$ from the first expert to perform region-scale analysis. By utilizing $\mathbf{\Phi}_{1}$ to mask irrelevant nodes in $\mathcal{G}$, it focuses on the emotion-relevant areas. The second expert then applies the same graph convolution and attention computation process with its own learnable parameters, producing a refined feature representation $\mathbf{H}_{2}$ and an attention map $\mathbf{\Phi}_{2}$ with important electrodes. The final stage of analysis is performed by the third expert, which combines $\mathcal{G}$ and $\mathbf{\Phi}_{2}$ to conduct electrode-scale analysis, yielding the feature representation $\mathbf{H}_{3}$ that captures fine-grained spatial patterns within EEG signals.